OpenStudy (anonymous):

(3x^7/2)^6(x^2)^6 What is the simplified form of the expression?

OpenStudy (anonymous):

( 3x - 6 ) / ( x - 2 ) = 3 * ( x - 2 ) / ( x - 2 ) = 3

OpenStudy (anonymous):

no the ^ mean squarout

OpenStudy (anonymous):

Oh my bad. In that case I really don't know. Sorry.